# references.bib

### v1

@@ Appendix A Simulation settings
As mentioned in Sec. 4.1, we create datasets of dynamic point clouds from simulations. All simulations are run with 200 points for 1,000 steps with a step size of $0.01$, and every 10th step is taken as an observation. Only for the final ablation experiment of Sec. 4.3 (i.e., Impact of observation timeframe), we simulate 20,000 time steps (of the D’Orsogna model) and extract training/testing sequences from this longer timeframe. The model parameters for these simulations are randomly sampled as specified below, and for each sampled parameter tuple, we create one simulation.
To create the dorsogna-10k dataset, we vary the following model parameters (see Fig. 3). Overall, we have four parameters that need to be predicted. As macroscopic regimes mainly depend on the ratios $C_{r}/C_{a}$ and $l_{r}/l_{a}$, cf. [Dorsogna06a, Fig. 1], we fix $C_{a}=l_{a}=1$ and sample $C_{r}=2^{t_{C}}$, $l_{r}=2^{t_{l}}$ with uniformly distributed $t_{C}\sim\mathcal{U}_{[-1,1]}$ and $t_{l}\sim\mathcal{U}_{[-1.5,0.5]}$. Similarly, we sample $\alpha=2^{t_{\alpha}}$ with $t_{\alpha}\sim\mathcal{U}_{[-2,2]}$ and $m=2^{t_{m}}$ with $t_{m}\sim\mathcal{U}_{[-2,2]}$. The model from the SiSyPHE library used to implement this simulation is AttractionRepulsion. Note, that the SiSyPHE library implements the mass $m$ in terms of the interaction radius parameter $R$, i.e., $m=R^{3}$ (as we simulate point clouds in 3D).
@@
For the vicsek-10k dataset, we sample $R,c,\nu$ uniformly from $\mathcal{U}_{[0.5,5.0]}$ and $D$ from $\mathcal{U}_{[0,2.0]}$. Overall, this gives four parameters that need to be predicted.
Finally, for the volex-10k dataset, we sample four parameters as follows: $\alpha\sim\mathcal{U}_{[0,2]}$, interaction radii $R\sim\mathcal{U}_{[0,2]}$ and birth/death rates $\lambda_{b},\lambda_{d}\sim\mathcal{U}_{[0,1]}$.
However, for the latter two, we discard settings where $\lambda_{d}>\lambda_{b}$, as in this case, death rates are almost impossible to distinguish, and also settings where $\lambda_{b}\gg\lambda_{d}$ in the sense that the resulting point cloud cardinalities exceed 2000 points during the simulation. The resulting $(\lambda_{b},\lambda_{d})$ combinations are illustrated in Fig. 5.
@@
Appendix B Comparison(s) to prior work
@@ B.1 Path Signature Kernel (PSK) [Giusti23a]
For the PSK, we rely on the publicly available implementation in sktime444 https://github.com/sktime/sktime [Loenig19a]. We compute the PSK (truncated at signature depth 3) using our vectorized persistence diagrams per time point as input. For concatenated zero-, one- and two-dimensional persistence diagrams, using 20-dimensional vectorizations (as specified in Sec. 4.2), this yields 60-dimensional input vectors per time point. The kernel is then input to a kernel support vector regressor (kernel-SVR).
Following the experimental protocol in [Giusti23a, Sec. 7.3], we cross-validate the sliding window embedding (across lags $[1,2,3]$) and kernel-SVR hyperparameters on a 20% validation portion of the training data, with hyperparameters selected based on the average MSE (per governing equation parameter to be predicted) across 5-folds. Due to the excessive runtime (approx. 8.3 hours per lag on the system specified in Appendix C), we only report results when all time points per observation sequence are considered. Hence, no subsampling of time points (as in our approach) is performed, and the reported performance can be considered an optimistic estimate of what can be achieved with the PSK.
@@ B.2 Crocker Stacks [Xian22a]
As mentioned in Sec. 4.4, we compare against crocker stacks, introduced in  [Xian22a], as one of the state-of-the-art approaches. Crocker stacks are an extension to crocker plots [Topaz15a] and constitute a topological summary for time-varying persistence diagrams. For the crocker stacks, we adapted the publicly available implementation of the crocker plots in teaspoon [^5] https://teaspoontda.github.io/teaspoon.
The crocker plot is computed as follows: for each time step, the persistence diagrams are computed up to a scale parameter $\varepsilon$. In discretized steps of the scale parameter, the Betti numbers are computed from the persistence diagrams, which results in a 2D representation for each homology dimension.
The extension to crocker stacks is achieved by adding a third dimension, induced by the smoothing factor $\alpha$. For given steps of $\alpha$, a smoothing operation is applied, i.e., values within a specified distance (smoothing factor) from the diagonal of the persistence diagram are ignored. Hence, for each homology dimension, a crocker stack is a tensor in $\mathbb{R}^{3}$, with axes corresponding to discretizations of the (1) scale parameter $\varepsilon\in[0,\infty)$, the (2) time $t\in[0,T]$, and the (3) smoothing factor $\alpha\in[0,\infty)$.
We note that in [Xian22a], the authors set the scale parameter to $\delta=0.35$, as they use data normalized to the range of 00 to $1$. We did not scale the data, however, we adjusted the scale parameter correspondingly: for all experiments, we set the scale parameter to $\delta=\nicefrac{{1}}{{3}}\cdot\text{maxPers}(\operatorname{dgm}_{k}(\text{Rips}(\mathcal{P})))$, where $\mathcal{P}$ denotes the point cloud and $\text{maxPers}(\operatorname{dgm}_{k}(\text{Rips}(\mathcal{P})))$ is the maximum persistence obtained from all point clouds in an observed sequence.
@@
Eventually, the crocker stacks per homology dimension are vectorized (i.e., the tensor is flattened) and concatenated into a single vector per observation sequence. These vectorizations are then input to a linear support vector regressor (SVR). For each of the (varied) parameters in the governing equation of Fig. 3, a separate SVR is trained and evaluated. Hyperparameters of the SVR and the usage of preprocessing steps (feature scaling) are tuned using Bayesian optimization. Each parameter configuration is evaluated using a 5-fold cross-validation and the best configuration is then used to train the final model on the full dataset.

### v2

@@ Appendix A Simulation settings
As mentioned in Sec. 4.1, we create datasets of dynamic point clouds from simulations. All simulations are run with 200 points for 1,000 steps with a step size of $0.01$, and every 10th step is taken as an observation. Only for the final ablation experiment of Sec. 4.3 (i.e., Impact of the observation timeframe), we simulate 20,000 time steps (of the D’Orsogna model) and extract training/testing sequences from this extended timeframe. The model parameters for these simulations are randomly sampled as specified below, and for each sampled parameter tuple, we create one simulation.
To create the dorsogna-10k dataset, we vary the following model parameters (see Fig. 3): overall, we have four parameters that need to be predicted. As macroscopic regimes mainly depend on the ratios $C_{r}/C_{a}$ and $l_{r}/l_{a}$, cf. [Dorsogna06a, Fig. 1], we fix $C_{a}=l_{a}=1$ and sample $C_{r}=2^{t_{C}}$, $l_{r}=2^{t_{l}}$ with uniformly distributed $t_{C}\sim\mathcal{U}_{[-1,1]}$ and $t_{l}\sim\mathcal{U}_{[-1.5,0.5]}$. Similarly, we sample $\alpha=2^{t_{\alpha}}$ with $t_{\alpha}\sim\mathcal{U}_{[-2,2]}$ and $m=2^{t_{m}}$ with $t_{m}\sim\mathcal{U}_{[-2,2]}$. The model from the SiSyPHE library used to implement this simulation is AttractionRepulsion. Note, that the SiSyPHE library implements the mass $m$ in terms of the interaction radius parameter $R$, i.e., $m=R^{3}$ (as we simulate point clouds in 3D).
@@
For the vicsek-10k dataset, we sample $R,c,\nu$ uniformly from $\mathcal{U}_{[0.5,5]}$ and $D$ from $\mathcal{U}_{[0,2]}$. Overall, this gives four parameters that need to be predicted.
Finally, for the volex-10k dataset, we sample four parameters as follows: $\alpha\sim\mathcal{U}_{[0,2]}$, interaction radii $R\sim\mathcal{U}_{[0,2]}$ and birth/death rates $\lambda_{b},\lambda_{d}\sim\mathcal{U}_{[0,1]}$. However, for the latter two parameters, we discard settings where $\lambda_{d}>\lambda_{b}$ as, in this case, death rates are almost impossible to distinguish; also, we discard settings where $\lambda_{b}\gg\lambda_{d}$ as the resulting point cloud cardinalities exceed 2000 points during the simulation. The resulting $(\lambda_{b},\lambda_{d})$ combinations are illustrated in Fig. 5.
@@
Appendix B Comparison(s) to prior work & Additional ablation experiments
@@ B.1 Path Signature Kernel (PSK) [Giusti23a]
For the PSK, we rely on the publicly available implementation in sktime444 https://github.com/sktime/sktime [Loenig19a]. We compute the PSK (truncated at signature depth 3) using our vectorized persistence diagrams per time point as input. For concatenated zero-, one- and two-dimensional persistence diagrams, we use 20-dimensional vectorizations (as specified in Sec. 4.2), yielding 60-dimensional input vectors per time point. The computed kernel is then input to a kernel support vector regressor (kernel-SVR).
Following the experimental protocol in [Giusti23a, Sec. 7.3], we cross-validate the sliding window embedding (across lags $[1,2,3]$) and kernel-SVR hyperparameters on a 20% validation portion of the training data, with hyperparameters selected based on the average MSE (per governing equation parameter $\beta_{i}$ to be predicted) across 5-folds. Due to the excessive runtime (approx. 8.3 hours per lag on the system specified in Appendix C), we only report results when all time points per observation sequence are considered. Hence, no subsampling of time points (as is done when evaluating our approach) is performed, and the reported performance can be considered an optimistic estimate of what can be achieved with the PSK.
@@ B.2 Crocker stacks [Xian22a]
As mentioned in Sec. 4.4, we compare against crocker stacks, introduced in [Xian22a], as one of the state-of-the-art approaches. Crocker stacks are an extension to crocker plots [Topaz15a] and constitute a topological summary for time-varying persistence diagrams. For the crocker stacks, we adapted the publicly available implementation of the crocker plots in the teaspoon [^5] https://teaspoontda.github.io/teaspoon library.
The crocker plot is computed as follows: for each time step, the persistence diagrams are computed up to a scale parameter $\varepsilon$. In discretized steps of the scale parameter, the Betti numbers are computed from the persistence diagrams, which results in a 2D representation (i.e., $\epsilon$ vs. Betti number) for each homology dimension.
The extension to crocker stacks is achieved by adding a third dimension $k$, induced by the smoothing factor $\alpha$. For given steps of $\alpha$, a smoothing operation is applied, i.e., values within a specified distance (smoothing factor) from the diagonal of the persistence diagram are ignored. Hence, for each homology dimension, a crocker stack is a tensor in $\mathbb{R}^{3}$, with axes corresponding to discretizations of the (1) scale parameter $\varepsilon\in[0,\infty)$, the (2) time $t\in[0,T]$, and the (3) smoothing factor $\alpha\in[0,\infty)$.
We note that in [Xian22a], the authors set the scale parameter to $\delta=0.35$, as they use data normalized to the range of 00 to $1$. We did not scale the data, however, we adjusted the scale parameter correspondingly: for all experiments, we set the scale parameter to $\delta=\nicefrac{{1}}{{3}}\cdot\text{maxPers}(\operatorname{dgm}_{k}(\text{Rips}(\mathcal{P})))$, where $\mathcal{P}$ denotes the point cloud, and $\text{maxPers}(\operatorname{dgm}_{k}(\text{Rips}(\mathcal{P})))$ is the maximum persistence obtained from all point clouds in an observed sequence.
@@
Eventually, the crocker stacks per homology dimension are vectorized (i.e., the tensor is flattened) and concatenated into a single vector per observation sequence. These vectorizations are then input to a linear support vector regressor (SVR). For each of the (varied) parameters in the governing equation of Fig. 3, a separate SVR is trained and evaluated. Hyperparameters of the SVR and the usage of preprocessing steps (feature scaling) are tuned using Bayesian optimization. Each parameter configuration is evaluated using a 5-fold cross-validation, and the best configuration is then used to train the final model on the full dataset.
B.3 Additional ablation experiments
To assess whether homology dimensions $>0$ are beneficial to the downstream regression task, we experimented on dorsogna-1k. We find that when using our approach with $\operatorname{H}_{0}$-features only, the regression quality drops. When additionally including $H_{2}$-features (i.e., $\operatorname{H}_{0}$, $\operatorname{H}_{1}$ and $\operatorname{H}_{2}$), the situation is less clear, as the results are not noticeably different. Quantitative results can be found in Tbl. 5 below.
Table 5: Quantitative assessment of the impact of including higher-dimensional homology features on downstream regression performance, evaluated on dorsogna-1k.
[…]
Ours (PH-only, v1; $\operatorname{H}_{0}$) 0.819 $\pm$ 0.015 0.101 $\pm$ 0.003
Ours (PH-only, v1; $\operatorname{H}_{0},\operatorname{H}_{1}$) 0.844 $\pm$ 0.021 0.098 $\pm$ 0.007
Ours (PH-only, v1; $\operatorname{H}_{0},\operatorname{H}_{1},\operatorname{H}_{2}$) 0.846 $\pm$ 0.011 0.097 $\pm$ 0.005
@@ Appendix C Computational resources
Appendix D Runtime analysis
In the following, we present a runtime breakdown of the pre-processing steps (i.e., Vietoris-Rips persistent homology (PH) computation, and the vectorization of persistence diagrams), as well as a runtime comparison to prior work (PSK and crocker stacks) and our baseline approach. Runtime is measured on the system listed above, using our publicly-available reference implementation.
At this point, it is also worth highlighting that our pre-processing step, i.e., PH computation & vectorization, is trivially parallelizable and can easily be distributed across multiple CPUs/GPUs if needed.
Vietoris-Rips PH computation. In Tbl. 6, we list wall clock time measurements (using Ripser++ [Zhang20a] on one GPU) per point cloud. In particular, the table lists runtime for computing PH of dimension zero and one (i.e., $\operatorname{H}_{0}$ and $\operatorname{H}_{1}$), as well as PH of dimension zero, one and two (i.e., $\operatorname{H}_{0},\operatorname{H}_{1}$ and $\operatorname{H}_{2}$). All point clouds in this experiment are of size 200 in $\mathbb{R}^{3}$.
Table 6: Runtime comparison for PH computation for different homology dimensions. Reported is the average runtime per point cloud (in seconds), and the overall runtime (in seconds) estimated from this average by multiplying by the number of processed point clouds on dorsogna-1k.
[…]
PH computation $(\operatorname{H}_{0},\operatorname{H}_{1})$ 0.018 s 1800 s
PH computation $(\operatorname{H}_{0},\operatorname{H}_{1},\operatorname{H}_{2})$ 0.330 s 33000 s
Persistence diagram vectorization. Persistence diagram (PD) vectorization can essentially be broken down into two steps: (1) parameter fitting for the structure elements of [Hofer19b] and (2) mapping PDs to vector representations using those structure elements. Tbl. 7 lists the runtime for both steps on the dorsogna-1k data when vectorizing zero-, and one-dimensional persistence diagrams. Throughout all of our experiments, parameter fitting for the structure elements is done by first collecting all persistence diagrams for each homology dimension and then running $k$-means++ clustering on 50,000 points uniformly sampled points from those diagrams.
Table 7: Runtime breakdown of PD vectorization (per diagram and overall for $\operatorname{H}_{0}$ and $\operatorname{H}_{1}$ on dorsogna-1k), split into (Step 1) time spent for fitting the parameters of the exponential structure elements from [Hofer19b], and (Step 2) actually mapping PDs to vector representations.
[…]
Step 1: Parameter fitting n/a 17 s
Step 2: Mapping PDs to vectors 9.4e-05 18 s
Comparison to prior work. Next, we compare the runtime of our approach (here, variant v1 from Fig. 2) to prior work that uses persistence diagrams as input. In particular, we compare against the PSK method from [Giusti23a] and crocker stacks [Xian22a].
Table 8: Training time comparison to prior work (on dorsogna-1k).
Training time
Crocker Stacks 24600 s
PSK 646 s
Ours (PH-only, v1) 190 s
In Tbl. 8, we list the overall training time (on dorsonga-1k), where runtime for pre-processing (see above) is excluded from these measurements. Also, PSK and crocker stacks timings do include hyperparameter optimization, as suggested in the corresponding references. Importantly, this is not optional but required to obtain decent performance with respect to EV and SMAPE. Notably, for the PSK approach, kernel computation scales quadratically with the number of sequences, and kernel-SVR training takes time somewhere between quadratic and cubic. Hence, scaling up the number of training sequences quickly becomes computationally prohibitive, especially in light of the required hyperparameter tuning. Finding suitable hyperparameters is also the main bottleneck for crocker stacks (which rely on a linear SVR).
Comparison to baseline model. Finally, in Tbl. 9, we present a training time comparison to our baseline model which does not explicitly model any dynamics via a latent ODE (denoted as w/o dynamics). In this experiment, the training protocol remains unchanged, and we vary the type of input data, i.e., from using vectorized persistence diagrams only ($\operatorname{H}_{0}$ and $\operatorname{H}_{1}$) to using a combination of PointNet++ features and vectorized persistence diagrams. Importantly, as remarked in the main part of the manuscript, the baseline models (PH-only, v1) and (PH+PointNet++, v3) already yields strong performance across all parameter prediction problems.
Table 9: Training time comparison (on dorsogna-1k) of our approach vs. the baseline model (w/o dynamics), using the same input data.
Training time
Ours (PH-only, v1) 190 s
Ours (PointNet++, v2) 3780 s
Ours (PH+PointNet++, v3) 4100 s
Baseline (w/o dynamics; PH-only, v1) 50 s
Baseline (w/o dynamics; PointNet++, v2) 525 s
Baseline (w/o dynamics; PH+PointNet++, v3) 600 s